Evaluate -1/8 - (-5/9)
Answer by ivy12003(22)

We can first simplify this question by combining the two negative signs together (negative and negative = positive). Then we find the common denominators by multiplying the two denominators together.



-1/8 - (-5/9) = -1/8 + 5/9 = (-1*9)/72 + (5*8)/72 = -9/72 + 40/72 = 31/72